If 𝑝 β†’ π‘ž is false, can you determine the truth value of (~𝑝)Λ…(𝑝 ↔ π‘ž)? Explain your answer.

If 𝑝 β†’ π‘ž is true, can you determine the truth value of ~(𝑝 β†’ π‘ž)Λ„~𝑝? Explain your answer.Β 


1
Expert's answer
2021-03-16T08:03:22-0400

If "p \to q \equiv F" then "p \equiv T,\,\,q \equiv F" . Then "\left( { \sim p} \right) \vee \left( {p \leftrightarrow q} \right) \equiv F \vee \left( {T \leftrightarrow F} \right) \equiv F \vee F \equiv F" . So, we can determine the truth value of (~𝑝)Λ…(𝑝 ↔ π‘ž).

f "p \to q \equiv T" then "\sim \left( {p \to q} \right) \equiv F" . But then "\sim \left( {p \to q} \right) \wedge \sim p \equiv F \wedge \sim p \equiv F" . So, we can determine the truth value of ~(𝑝 β†’ π‘ž)Λ„~𝑝.